OpenStudy (joey):

judith went shopping with $50 and spent $45 what is the percentage of money she spent 90% 80% 95% 85%

jhonyy9 (jhonyy9):

use the rule of 3 simple $50 mean 100% $45 ---- x ------------------ x = 45*100/50 = 45*2 = 90% so this is the right answer - hope you understand it now was my pleasure